# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 6.2 Alternating direction method of multipliers (ADMM)
Due to the equivalence between ADMM and DRS, we now transform the ADMM into a DRS form. Let $z^{k}=u^{k}+\tau B^{\top}t^{k}$, then, (17) becomes:
$\left\lfloor\begin{array}[]{lll}t^{k}&\in&\operatorname{Arg}\min_{t}g(t)+\frac{\tau}{2}\big{\|}Bt-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ s^{k+1}&\in&\operatorname{Arg}\min_{s}f(s)+\frac{\tau}{2}\big{\|}As+2Bt^{k}-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ z^{k+1}&=&z^{k}-\tau(As^{k+1}+Bt^{k}).\\ \end{array}\right.$ (18)
Similar to [33, Remark 3.1-(ii)], (18) can be rewritten in terms of infinal postcomposition [4, Definition 12.34]:
@@
The limiting case of the well-known Chambolle–Pock (CP) algorithm [14, Algorithm 1] can also be analyzed in a similar manner, based an augmentation technique [25], which equivalently expresses the CP algorithm as the DRS form (also see [33, Sect. 5.3] for details).
@@ 6.3 Augmented Lagrangian method (ALM)
where $B:\mathcal{U}\mapsto\mathcal{Y}$. The ALM reads as (see [31, Eq.(7.2)] for example):
@@
Using the notion of infimal postcomposition mentioned above, we obtain the equivalent form to (20):
@@
First, observing (22) obtains that $\mathcal{A}$ is maximally monotone, $\mathsf{ran}\mathcal{A}=\mathcal{Y}^{2}=\mathcal{H}$, and thus, $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$, which fulfills Assumption 4. Then it follows that the scheme (21) admits any input $x=(q,p)\in\mathcal{Y}^{2}$.
Then, we check Theorem 4.2 and 6. Here, $\mathcal{E}=\begin{bmatrix}I&0\end{bmatrix}$. Given an input $x_{\mathsf{r}}=(0,a)$ and the corresponding output $y_{\mathsf{r}}=(0,c)$, the inclusion of $\mathcal{A}(\mathcal{E}^{\top}v+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ becomes:
@@
The second line yields that $v=\frac{1}{\tau}(a-c)+b$, which is obviously single-valued (satisfying Theorem 4.2). And furthermore, $\|v-v^{\prime}\|=\frac{1}{\tau}\|(a-a^{\prime})-(c-c^{\prime})\|\leq\frac{2}{\tau}\|a-a^{\prime}\|$, i.e., $\frac{2}{\tau}$-Lipschitz continuous w.r.t. $a$. Finally, the weak convergence of $\{q^{k},p^{k}\}_{k\in\mathbb{N}}$ follows by Theorem 5.6.
@@ 8 Acknowledgements
We are gratefully indebted to Mr. Brecht Evens (KU Leuven, Belgium) for inspiring comments and bringing [22, Theorem 3.4] to our attention in the early stages of this work. We would also like to thank Mr. Shuchang Zhang (NUDT, China) for detailed discussions on [33, Theorem 2.1].
@@
Lemma A.1 extends the classical Moreau’s identity to the degenerate case. This is a key ingredient for proving Lemma 2.10. We stress that this extension holds under Assumption 1 only, even without the restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ only. This greatly weakens the previous result of [33, Lemma 2.3].
@@
Under Assumption 1, we have $\mathcal{I}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}+(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$, for any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$;
@@
For any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$, pick $y$, such that $y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we develop $\mathcal{Q}x\in\mathcal{A}y+\mathcal{Q}y\Longrightarrow\mathcal{Q}(x-y)\in\mathcal{A}y\Longrightarrow y\in\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in x-\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Conversely, For any input $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, pick $y$, such that $y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$. Then, we develop $y+\mathcal{A}^{-1}\mathcal{Q}y\owns x\Longrightarrow x-y\in\mathcal{A}^{-1}\mathcal{Q}y\Longrightarrow\mathcal{Q}y\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}(y-x+x)\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-y)\Longrightarrow x-y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\supseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Both directions show the desired result, with $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Let us now show that all the expressions appear in Lemma 2.10 are single-valued under Assumptions 1 and 2.
@@
(i) Given $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, let $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$, then, $y_{1}+\mathcal{A}^{-1}\mathcal{Q}y_{1}\owns x$ and $y_{2}+\mathcal{A}^{-1}\mathcal{Q}y_{2}\owns x$, leading to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$. Then,
@@
where the last inequality is due to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$ and the local monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Finally, we have $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, which shows the single-valuedness of $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$.
(ii) Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y_{1}+\mathcal{Q}\mathcal{A}^{-1}y_{1}\owns\mathcal{Q}x$ and $y_{2}+\mathcal{Q}\mathcal{A}^{-1}y_{2}\owns\mathcal{Q}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathsf{ran}\mathcal{Q}$. Let $y_{1}=\mathcal{Q}v_{1}$ and $y_{2}=\mathcal{Q}v_{2}$, we have $\mathcal{Q}v_{1}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\mathcal{Q}x$ and $\mathcal{Q}v_{2}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\mathcal{Q}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}$, we have:
@@
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$}$
@@
(iii) It suffices to show that $\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}$ is single-valued. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$. Then, $y_{1}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y_{1}\owns\sqrt{\mathcal{Q}}x$ and $y_{2}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y_{2}\owns\sqrt{\mathcal{Q}}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathsf{ran}\mathcal{Q}$. Let $y_{1}=\sqrt{\mathcal{Q}}v_{1}$ and $y_{2}=\sqrt{\mathcal{Q}}v_{2}$, we have $\sqrt{\mathcal{Q}}v_{1}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\sqrt{\mathcal{Q}}x$ and $\sqrt{\mathcal{Q}}v_{2}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\sqrt{\mathcal{Q}}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$, we have:
$=$ $\langle v_{1}-v_{2}|\mathcal{Q}(s_{1}-v_{1}-s_{2}+v_{2})\rangle$
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$}$
which leads to $\mathcal{Q}v_{1}=\mathcal{Q}v_{2}$, i.e., $y_{1}=y_{2}$.
@@
(2): We begin with $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$. Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$. Let $v\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$ and $y=\mathcal{Q}v$. Then, we deduce that $v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{Q}v+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v\owns\mathcal{Q}x\Longrightarrow y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x\Longrightarrow y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$. This also shows that $\mathsf{dom}\big{(}\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$.
(3): We begin with $(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}$. Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Let $y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$. Let $y=\sqrt{\mathcal{Q}}v$ with $v\in\mathsf{ran}\mathcal{Q}$ [^4] Strictly speaking, we should write $y=\sqrt{\mathcal{Q}}(v+s)$ with $s\in\mathsf{ker}\mathcal{Q}$. This is essentially equivalent to $y=\sqrt{\mathcal{Q}}v$.. Then, the inclusion becomes $\sqrt{\mathcal{Q}}v+\mathcal{Q}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v\owns\mathcal{Q}x\Longrightarrow v+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v\owns\sqrt{\mathcal{Q}}x\Longrightarrow v=\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x\Longrightarrow y=\sqrt{\mathcal{Q}}v=\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$. This also shows that $\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}\subseteq\mathsf{dom}\big{(}\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$.
(4): It suffices to show $\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}$. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Let $y=\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$, then, $y+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y\owns\sqrt{\mathcal{Q}}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$, which already satisfies the domain restricted in $\mathsf{ran}\mathcal{Q}$, i.e., $y+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}|_{\mathsf{ran}\mathcal{Q}}(y)\owns\sqrt{\mathcal{Q}}x$, leading to the right-hand side of (4).
(4$\Longrightarrow$(1): It suffices to show that $\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}=\sqrt{\mathcal{Q}}-\sqrt{\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Let $y=\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$, then, $y+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}(y)\owns\sqrt{\mathcal{Q}}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$, and thus, the above inclusion can be simplified as $y+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y\owns\sqrt{\mathcal{Q}}x$, and let $y=\sqrt{\mathcal{Q}}v$ with $v\in\mathsf{ran}\mathcal{Q}$. Then, we have $\sqrt{\mathcal{Q}}v+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v\owns\sqrt{\mathcal{Q}}x\Longrightarrow v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{A}^{-1}\mathcal{Q}v\owns x-v\Longrightarrow\mathcal{Q}v\in\mathcal{A}(x-v)\Longrightarrow\mathcal{Q}(v-x+x)\in\mathcal{A}(x-v)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-v)\Longrightarrow x-v\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow v\in x-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow y=\sqrt{\mathcal{Q}}v=\sqrt{\mathcal{Q}}(\mathcal{I}-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q})x$. This also shows that $\mathsf{dom}\big{(}\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$.
Finally, we show all the equalities with exactly same domains.

### v4

@@ 6.2 Alternating direction method of multipliers (ADMM)
Similar to [51, Remark 3.1-(ii)], letting $z^{k}=u^{k}+\tau B^{\top}t^{k}$, and using infinal postcomposition [5, Definition 12.34], (17) becomes:
@@
The limiting case of the well-known Chambolle–Pock (CP) algorithm [20, Algorithm 1] can also be analyzed in a similar manner, based an augmentation technique [39], which equivalently expresses the CP algorithm as the DRS form (also see [51, Sect. 5.3] for details).
@@ 6.3 Augmented Lagrangian method (ALM)
where $B:\mathcal{U}\mapsto\mathcal{Y}$. The ALM reads as (see [48, Eq.(7.2)] for example):
@@
This is equivalent to:
@@
First, observing (21) obtains that $\mathcal{A}$ is maximally monotone, $\mathrm{ran}\mathcal{A}=\mathcal{Y}^{2}=\mathcal{H}$, and thus, $0\in\mathrm{sri}(\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A})$, which fulfills Assumption 4. Then it follows that the scheme (20) admits any input $x=(q,p)\in\mathcal{Y}^{2}$.
Then, we check Theorem 4.2 and 5. Here, $\mathcal{E}=\begin{bmatrix}I&0\end{bmatrix}$. Given an input $x_{\mathsf{r}}=(0,a)$ and the corresponding output $y_{\mathsf{r}}=(0,c)$, the inclusion of $\mathcal{A}(\mathcal{E}^{\top}v+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ becomes:
@@
The second line yields that $v=\frac{1}{\tau}(a-c)+b$, which is obviously single-valued (satisfying Theorem 4.2). Furthermore, $\|v-v^{\prime}\|=\frac{1}{\tau}\|(a-a^{\prime})-(c-c^{\prime})\|\leq\frac{2}{\tau}\|a-a^{\prime}\|$, i.e., $\frac{2}{\tau}$-Lipschitz continuous w.r.t. $a$. Finally, the weak convergence of $\{q^{k},p^{k}\}_{k\in\mathbb{N}}$ follows by Theorem 5.6.
@@ 8 Acknowledgements
This work was supported by the National Natural Science Foundation of China under grant 12471300. We are gratefully indebted to Mr. Brecht Evens (KU Leuven, Belgium) for inspiring comments and bringing [33, Theorem 3.4] to our attention in the early stages of this work. We would also like to thank Prof. Jianchao Bai (NWPU, China) and Mr. Shuchang Zhang (NUDT, China) for detailed discussions on ADMM algorithms (17) and [51, Theorem 2.1].
@@
Lemma A.1 extends the classical Moreau’s identity to the degenerate case. This is a key ingredient for proving Lemma 2.11. We stress that this extension holds under Assumption 1 only, even without the restricted monotonicity of $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ only. This greatly weakens the previous result of [51, Lemma 2.3].
@@
Under Assumption 1, we have $\mathcal{I}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}+(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$, for any input $x\in\mathrm{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$;
@@
For any input $x\in\mathrm{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$, pick $y$, such that $y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we develop $\mathcal{Q}x\in\mathcal{A}y+\mathcal{Q}y\Longrightarrow\mathcal{Q}(x-y)\in\mathcal{A}y\Longrightarrow y\in\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in x-\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathrm{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\subseteq\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Conversely, for any input $x\in\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, pick $y$, such that $y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$. Then, we develop $y+\mathcal{A}^{-1}\mathcal{Q}y\owns x\Longrightarrow x-y\in\mathcal{A}^{-1}\mathcal{Q}y\Longrightarrow\mathcal{Q}y\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}(y-x+x)\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-y)\Longrightarrow x-y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathrm{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\supseteq\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Both directions show the desired result, with $\mathrm{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Let us now show that all the expressions (composite operators) appear in Lemma 2.11 are single-valued under Assumptions 1 and 2.
@@
(i) Given $x\in\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, let $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$, then, $y_{1}+\mathcal{A}^{-1}\mathcal{Q}y_{1}\owns x$ and $y_{2}+\mathcal{A}^{-1}\mathcal{Q}y_{2}\owns x$, leading to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$. Then,
@@
where the last inequality is due to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$ and the local monotonicity of $\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$. Finally, we have $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, which shows the single-valuedness of $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$.
(ii) Take $x\in\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y_{1}+\mathcal{Q}\mathcal{A}^{-1}y_{1}\owns\mathcal{Q}x$ and $y_{2}+\mathcal{Q}\mathcal{A}^{-1}y_{2}\owns\mathcal{Q}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathrm{ran}\mathcal{Q}$. Let $y_{1}=\mathcal{Q}v_{1}$ and $y_{2}=\mathcal{Q}v_{2}$, we have $\mathcal{Q}v_{1}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\mathcal{Q}x$ and $\mathcal{Q}v_{2}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\mathcal{Q}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathrm{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}$, we have:
@@
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathrm{ker}\mathcal{Q}$}$
@@
(iii) Similar to (ii), omitted.
@@
(2): We begin with $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$. Take $x\in\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$. Let $v\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$ and $y=\mathcal{Q}v$. Then, we deduce that $v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{Q}v+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v\owns\mathcal{Q}x\Longrightarrow y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x\Longrightarrow y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$. This also shows that $\mathrm{dom}\big{(}\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}\subseteq\mathrm{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$.
The deductions of (3$\Longrightarrow$(4) and (4$\Longrightarrow$(1) are similar to the proof of (2$\Longrightarrow$(3), and thus omitted here.